OpenStudy (anonymous):

Two cylinders are similar with a surface area ratio of 25:64. Find the volume of the second given that the volume of the first is 250 m3. Answer A. 61.0 m3 B. 640 m3 C. 1024 m3 D. 97.7 m3

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

hello lad are you willying to help

OpenStudy (anonymous):

*willing

OpenStudy (anonymous):

@gitahimart are you going to help :):)(:(:

OpenStudy (anonymous):

do i set it up as a cross multiply thing

OpenStudy (anonymous):

Help kind sir please i beg of you

OpenStudy (anonymous):

@Liltico sorry was out, consider this: Use these two facts about similar objects. Area ratio = (length ratio)^2 Volume ratio = (length ratio)^3 Find the length ratio by square rooting the area ratio and then use it in the second formula., can you do that?

OpenStudy (anonymous):

you deserve a metal.

OpenStudy (anonymous):

thanks, what have you got?

OpenStudy (anonymous):

sir please be more detailed i don't understand

OpenStudy (anonymous):

OK, you,ve been given an area (surface area) ration of 25/64, so the length ratio is the square root of this which is 5/8, following upto there?

OpenStudy (anonymous):

I got B am i correct

OpenStudy (anonymous):

nope, what was your length ratio?

OpenStudy (anonymous):

5/64

OpenStudy (anonymous):

remember: \[\sqrt{(25/64)} = (\sqrt{25})/(\sqrt{64}) = 5/8\]

OpenStudy (anonymous):

now from here you need to get the volume ratio, and since volume means cubing values, then the volume ratio becomes: \[5^{3}:8^{3} = 125:512\]

OpenStudy (anonymous):

@Liltico

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

we together?

OpenStudy (anonymous):

i understand how you got 125:512 but i'm lost in finding the answers

OpenStudy (anonymous):

Don't you leave me!

OpenStudy (anonymous):

we're getting there,now since you've been told that the volume of the 1st is 250, from the ratio you can see that the volume of the 2nd will be larger..

OpenStudy (anonymous):

so: 125:512 250:x

OpenStudy (anonymous):

X=61

OpenStudy (anonymous):

Is that my answer?

OpenStudy (anonymous):

to find x, cross multiply: x * 125 = 250 * 512 x = (250*512)/125 x =1024

OpenStudy (anonymous):

1024 is your answer

OpenStudy (anonymous):

oh i understand thanks once more

OpenStudy (anonymous):

welcome